Question 2. (24 points. Recall that in Assignment 1, we developed recursive algorithms for implementing the specification GivQuestion 2(A). (4p) Instantiate AA to assignments that establish (you do not need to argue why).
Answer #1

Determine AA on the basis of given \Phi

Add statements to AA which initializes the conditions for while loop using the 3 given conditions.

1. 1 \le lo and hi \le n

Initially lo will at the start index and hi will be at the end index.

=> lo <- 1, hi <- n

2. If v occurs in A[1...n] then v occurs in A[lo...hi]

This condition also determines the above inferred information for AA, that the range of index for A is lo to hi.

And initiallly values for lo is 1 and for hi is n.

3. If the boolean variable found is true then 1 \le q \le n and A[q] = v

If found becomes true, then v is found in A array. Therefore, initially initialize found as false, so that as soon as v is encountered in A, found becomes true.

=> found <- false

Hence putting all initial conditions together,

AA = lo <- 1, hi <- n, found <- false